The joint pdf ofX and Y is f (x,y)=kxy







1
Expert's answer
2021-11-29T13:36:11-0500

for "0<x<1,0<y<1"


"\int^1_0\int^1_0f(x,y)dxdy=1"


"\int^1_0\int^1_0kxydxdy=k\int^1_0(y/2)dy=k/4=1"


"k=4"